OpenStudy (anonymous):

find the inverse of f(x)=1/x-2

OpenStudy (anonymous):

f(x)=1/x -2 ==> f(x)+2 = 1/x ==> x= 1/ ( f(x)+2)

OpenStudy (anonymous):

inverse function 1/x-2 result: 1/(x+2)

Directrix (directrix):

@cutegirl Is this f(x)=1/x-2 meant to be f(x) = (1/x) -2 OR f(x) = 1/(x-2) ?